Consider the following data for a dependent variable and two independent variables x1 and x2 .

x1 x2 y
30 13 94
46 11 109
25 17 113
51 16 179
40 5 95
52 19 175
74 7 171
37 13 118
59 14 142
76 17 211

Round your all answers to two decimal places. Enter negative values as negative numbers, if necessary.

a. Develop an estimated regression equation relating y to x1 .

Predict y if x1=35 .

y-hat=

b. Develop an estimated regression equation relating y to x2 .

Predict y if x2=15 .

y-hat=

c. Develop an estimated regression equation relating y to x1 and x2 .

Predict y if x1=35 and x2=15 .

y-hat=
Answer #1

a.

Estimated equation:

Y = 44.50 + 1.96 X1

When X1 = 35,

Y-hat = 44.50 + 1.96(35) = 113.10

b.

Estimated equation:

Y = 86.94 + 4.07 X2

When X2 = 15,

Y-hat = 86.94 + 4.07(15) = 147.99

c.

Estimated equation:

Y = -15.15 + 2.00 X1 + 4.36 X2

When X1=35 and X2=15

Y-hat = -15.15 + 2.00 (35) + 4.36 (15) = 120.25

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!